Calculate the mass of oxalic acid needed to make a 250.0 mL standard solution that is 0.0500 mol L-1 . You will use oxalic acid dihydrate, (COOH)2.2H2O. M((COOH)2.2H2O) = 126.07 g mol-1 )
Answer #1

Solution-

suppose w gram oxalic acid needed.

Volume = 250ml =250× 10^-3L=0.25L

Molar mass = 126.07 gram/mole

Moles =w/126.07 mole

Molarity = moles/volume

=w/126.07×0.25

0.0500=w/31.52

w=0.0500×31.52

w=1.58 gram.

Hence, 1.58 gram required.